Kelly brought 8 apples, 3 pears, and 13 oranges and brought them home in a grocery bag. If she picks one from the bag without looking what is the probability of choosing an apple? a pear? an orange?
The chances of choosing an apple are________?
The chances of choosing an pear are_________?
The chances of choosing an orange are_______?

Answer by stanbon(75887)

Kelly brought 8 apples, 3 pears, and 13 oranges and brought them home in a grocery bag. If she picks one from the bag without looking what is the probability of choosing an apple? a pear? an orange?
--------------------------------------------------------------
The chances of choosing an apple are 8/24 = 1/3
The chances of choosing an pear are 3/24 = 1/8
The chances of choosing an orange are 13/24
==============================================
Cheers,
Stan H.